# Investigating a Model-Agnostic and Imputation-Free Approach for Irregularly-Sampled Multivariate Time-Series Modeling

### v2

@@ A.2 Classification
Epilepsy. Epilepsy (Andrzejak et al. 2001) contains univariate brainwaves (single-channel EEG) sampled from 500 subjects (with 11,500 samples in total), with each sample classified as having epilepsy or not (binary classification).
Gesture. Gesture (Liu et al. 2009) dataset consists of 560 samples, each having 3 variates (corresponding to the accelerometer data) and each sample corresponding to one of the 8 hand gestures (or classes)
EMG. EMG (Goldberger et al. 2000a) dataset contains 163 EMG (Electromyography) samples corresponding to 3-classes of muscular diseases.
@@
Physio-Net Dataset: PhysioNet-2012 Mortality Prediction Challenge (Silva et al. 2012) contains 12k multivariate clinical time-series samples that were collected from patients in ICU. The time-series contains 37 variables, such as temperature, heart rate, blood pressure, etc. that can vary depending on the type of patient. Each of the samples are recorded during the first 48 hours of admission in ICU. PhysioNet has a high degree of 80% missing values. We follow the experimental setup in (Du, Côté, and Liu 2023), and split the dataset into 80%, 10% and 10% train/val/test split.
P12 Dataset Derived from the PhysioNet 2012 Challenge, the P12 (Goldberger et al. 2000b) dataset includes multivariate ICU time-series data for 11,988 patients after filtering out 12 invalid entries. Each patient record contains up to 48 hours of sensor data across 36 physiological variables (excluding weight), plus a 9-dimensional static vector with demographic attributes such as age and gender. The binary label indicates ICU length of stay: $\leq$3 days (negative) vs. $\geq$3 days (positive).
P19 Dataset (Reyna et al. 2020) This dataset is part of the PhysioNet 2019 Sepsis Early Prediction Challenge and contains longitudinal EHR data from 38,803 ICU patients. Each patient record comprises multivariate time-series data with 34 clinical variables, sampled at irregular intervals, alongside a static vector containing demographic and admission information (e.g., age, gender, ICU type, hospital-to-ICU delay, ICU length of stay). A binary label indicates whether the patient will develop sepsis within the next 6 hours.
We report results on P12 and P19 dataset in the appendix only, as supplementary evaluation for IMTS classification.
@@ Appendix B Baselines
Regular MTS Methods. iTransformer (Liu et al. 2023), PatchTST (Nie et al. 2022b), DLinear (Zeng et al. 2023), Autoformer (Wu et al. 2021), SimMTM (Dong et al. 2024), MAE (He et al. 2022) (adapted for Time series similar to Ti-MAE (Li et al. 2023)). We use the default parameters reported for each models in their official code for the respective datasets. We use the parameters reported in Table 4 and Table 6 for the MAE implementation, as code for Ti-MAE (Li et al. 2023) is not publicly available.
Imputation Methods $2^{\text{nd}}$-order spline imputation (McKinley and Levine 1998), k-Nearest Neighbor (kNN) (Tan et al. 2019) (with neighbors=10), SAITS (Du, Côté, and Liu 2023), and BRITS (Cao et al. 2018). For SAITS, we use the following setting in our experiments n_layers=2, embedding_dim=256, num_heads=4. For BRITS, we use hidden_dimension=64. We follow the official code for implementation of these models.
IMTS Methods GRU-D (Che et al. 2018), Latent ODE (Rubanova, Chen, and Duvenaud 2019). We used the available GRU-D re-implementation as the official code was not available. For Latent ODE we used their official code. The additional experiments reported in the appendix include the following baselines, SeFT (Horn et al. 2020), mTAND (Shukla and Marlin 2021), Raindrop (Zhang et al. 2021), MTGNN (Wu et al. 2020)
@@ Appendix C Implementation Details
The experiments have been implemented in PyTorch using NVIDIA TITAN 24 GB GPU. The baselines have been implemented following their official code and configurations. We consider Mean Squared Error (MSE) as the metric for time-series forecasting and F1-score for the classification tasks. We generate five different versions of synthetic data and report the average metric from these 5 runs along with their standard deviation, wherever applicable. We report AUROC and AUPRC for some additional experiments included in the supplementary.
@@ C.1 Hyper-parameter Details
The hyperparameters for MissTSM are selected after integration with the base model, with embedding dimensions searched in the range [4, 128] and the number of attention heads in the range [1, 8]. For the base models, we use the default configurations for the iTransformer (Liu et al. 2023) and PatchTST (Nie et al. 2022b) models. For the MAE base model, we start with the same set of hyper-parameters as reported in the SimMTM paper (see Table 4) as initialization, and then search for the best learning rate in factors of 10, encoder/decoder layers in the range [2, 4], number of heads in the range [2,16] and embedding dimensions in the range [4, 128]. Note that we only perform hyper-parameter tuning on 100% (or fully observed) data, and use the same hyper-parameters for all experiments involving the dataset, such as different missing value probabilities. Table 5 captures the MissTSM parameters across datasets. Note that, for classification datasets (EMG, Epilepsy, Gesture), we use the same set of parameters with $q_{\text{dim}}=v_{\text{dim}}=k_{\text{dim}}=32$ and $n_{\text{head}}=16$. Table 6 presents the sensitivity analysis of the MissTSM-MAE model with respect to key hyperparameters (measured using MSE).
@@ C.2 Forecasting experiments
The models were trained with the MSE loss, using the Adam (Kingma 2014) optimizer with a learning rate of 1e-3 during pre-training for 50 epochs and a learning r ate of 1e-4 during finetuning with an early stopping counter of 3 epochs. Batch size was set to 16. All the reported missing data experiment results are obtained over 5 trials (5 different masked versions). During fine-tuning for different Prediction lengths (96, 192, 336, 720), we used the same pre-trained encoder and added a linear layer at the top of the encoder.
@@ C.3 Classification experiments
The models were trained using the Adam (Kingma 2014) optimizer, with MSE as the loss function during pre-training and Cross-Entropy loss during fine-tuning. During fine-tuning, we plugged a 64-D linear layer at the top of the pre-trained encoder. We pre-trained and fine-tuned for 100 epochs. For the GRU-D experiments, we use the available re-implementation of the code, with hidden size as 16. For the PhysioNet experiments, we follow the same evaluation setup as proposed in (Du, Côté, and Liu 2023). The baselines imputation models use a simple RNN model as a classification model on top of the imputed data.
Appendix D Computational complexity and error analysis
@@
We conduct a case study to examine the trade-offs regular MTS models for IMTS modeling in terms of computational complexity and efficiency, and to investigate potential correlations between imputation errors and downstream performance. For that, we consider a classification task on the Epilepsy dataset. Dataset is 80% masked under MCAR. Spline and SAITS are the imputation techniques and SimMTM is the MTS model used. We report the total modeling time as the sum of imputation time and the time-series model training time. The experiments are conducted on NVIDIA TITAN 24 GB GPUs.
@@ E.3 Additional Forecasting Dataset
To further strengthen our understanding of the behaviors of regular IMTS models under the two-step pipeline, and to explore the potential of model-agnostic imputation-free approaches, we conduct forecasting experiments on another dataset, Solar-Energy (Lai et al. 2018), using the MCAR masking scheme over two missing data fractions - 60 and 70%. Figure 10 summarizes the results. As expected, the performance of traditional MTS methods degrades with increasing missingness, reflecting their reliance on separate imputation strategies. While IMTS methods provide a practical solution for partially observed time series, they tend to underperform compared to regular MTS models in long-range forecasting scenarios. This is where we believe the gap could potentially be addressed by imputation-free, model-agnostic approaches such as MissTSM.
Figure 10: Results on the Solar-Energy dataset with MCAR masking across four horizons at 60% and 70% masking fractions.
E.4 IMTS Classification
We conduct additional experiments on the IMTS classification task using the P12 (Goldberger et al. 2000b) and P19 (Reyna et al. 2020) datasets, following the same evaluation protocol as (Luo et al. 2025). We report the baseline results for the considered models directly from (Luo et al. 2025). Table 10 highlights the strong potential of model-agnostic approaches; integrating the MissTSM layer, can achieve performance on par with or exceeding that of several well-known IMTS models.
Methods P19 P12
AUROC AUPRC AUROC AUPRC
GRU-D 88.7 1.2 56.2 2.3 79.6 0.6 41.7 1.8
ODE-RNN 87.1 1.0 52.6 3.2 78.8 0.6 37.4 2.6
SeFT 84.0 0.3 49.3 0.5 78.1 0.5 35.9 0.8
mTAND 82.9 0.9 32.2 1.5 85.3 0.3 49.3 1.0
Raindrop 87.6 2.7 61.1 1.4 82.0 0.6 42.7 1.7
MTGNN 88.5 1.0 55.8 1.5 82.1 1.5 41.8 2.1
MissTSM 88.8 1.3 56.5 1.2 82.2 0.5 43.8 1.1
E.5 Ablations on forecasting & classification tasks
In the ablation experiments, our goal is to quantify the effectiveness of plugging in the MissTSM layer. To achieve this, we compare MAE with MissTSM vs MAE without MissTSM, which we refer to as just MAE. For both the forecasting (see Fig. 11) and classification (see Fig. 12) tasks, we compare the model integrated with MissTSM against MAE trained on spline and SAITS imputation techniques. Under the forecasting task, we observe that plugging in the MissTSM layer consistently outperforms the MAE ablations. Under classification tasks, for all the three datasets, we observe that we MissTSM-MAE is either comparable or better than the MAE ablation.
E.6 Analyzing the Impact of Forecast Horizon

### v3

@@ A.2 Classification
Epilepsy. Epilepsy (Andrzejak et al., 2001) contains univariate brainwaves (single-channel EEG) sampled from 500 subjects (with 11,500 samples in total), with each sample classified as having epilepsy or not (binary classification).
Gesture. Gesture (Liu et al., 2009) dataset consists of 560 samples, each having 3 variates (corresponding to the accelerometer data) and each sample corresponding to one of the 8 hand gestures (or classes)
EMG. EMG (Goldberger et al., 2000a) dataset contains 163 EMG (Electromyography) samples corresponding to 3-classes of muscular diseases.
@@
Physio-Net Dataset: PhysioNet-2012 Mortality Prediction Challenge (Silva et al., 2012) contains 12k multivariate clinical time-series samples that were collected from patients in ICU. The time-series contains 37 variables, such as temperature, heart rate, blood pressure, etc. that can vary depending on the type of patient. Each of the samples are recorded during the first 48 hours of admission in ICU. PhysioNet has a high degree of 80% missing values. We follow the experimental setup in Du et al. (2023), and split the dataset into 80%, 10% and 10% train/val/test split.
P12 Dataset Derived from the PhysioNet 2012 Challenge, the P12 (Goldberger et al., 2000b) dataset includes multivariate ICU time-series data for 11,988 patients after filtering out 12 invalid entries. Each patient record contains up to 48 hours of sensor data across 36 physiological variables (excluding weight), plus a 9-dimensional static vector with demographic attributes such as age and gender. The binary label indicates ICU length of stay: $\leq$3 days (negative) vs. $\geq$3 days (positive).
P19 Dataset (Reyna et al., 2020) This dataset is part of the PhysioNet 2019 Sepsis Early Prediction Challenge and contains longitudinal EHR data from 38,803 ICU patients. Each patient record comprises multivariate time-series data with 34 clinical variables, sampled at irregular intervals, alongside a static vector containing demographic and admission information (e.g., age, gender, ICU type, hospital-to-ICU delay, ICU length of stay). A binary label indicates whether the patient will develop sepsis within the next 6 hours.
@@ Appendix B Baselines
Regular MTS Methods. iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022b), DLinear (Zeng et al., 2023), Autoformer (Wu et al., 2021), SimMTM (Dong et al., 2024), MAE (He et al., 2022) (adapted for Time series similar to Ti-MAE (Li et al., 2023)). We use the default parameters reported for each models in their official code for the respective datasets. We use the parameters reported in Table 6 and Table 8 for the MAE implementation, as code for Ti-MAE (Li et al., 2023) is not publicly available.
Imputation Methods $2^{\text{nd}}$-order spline imputation (McKinley and Levine, 1998), k-Nearest Neighbor (kNN) (Tan et al., 2019) (with neighbors=10), SAITS (Du et al., 2023), and BRITS (Cao et al., 2018). For SAITS, we use the following setting in our experiments n_layers=2, embedding_dim=256, num_heads=4. For BRITS, we use hidden_dimension=64. We follow the official code for implementation of these models.
IMTS Methods GRU-D (Che et al., 2018), Latent ODE (Rubanova et al., 2019). We used the available GRU-D re-implementation as the official code was not available. For Latent ODE we used their official code. The other baselines used are the following, SeFT (Horn et al., 2020), mTAND (Shukla and Marlin, 2021), Raindrop (Zhang et al., 2021), MTGNN (Wu et al., 2020)
@@ Appendix C Implementation Details
The experiments have been implemented in PyTorch using NVIDIA TITAN 24 GB GPU. The baselines have been implemented following their official code and configurations. In the default implementation code of MissTSM with MAE, we integrate the MissTSM layer directly within the backbone without the linear projection to the original input shape. We consider Mean Squared Error (MSE) as the metric for time-series forecasting and F1-score, AUROC, AUPRC for the classification tasks. We generate five different versions of synthetic data and report the average metric from these 5 runs along with their standard deviation, wherever applicable.
@@ C.1 Hyper-parameter Details
The hyperparameters for MissTSM are selected after integration with the base model, with embedding dimensions searched in the range [4, 128] and the number of attention heads in the range [1, 8]. For the base models, we use the default configurations for the iTransformer (Liu et al., 2023) and PatchTST (Nie et al., 2022b) models. For the MAE base model, we start with the same set of hyper-parameters as reported in the SimMTM paper (see Table 6) as initialization, and then search for the best learning rate in factors of 10, encoder/decoder layers in the range [2, 4], number of heads in the range [2,16] and embedding dimensions in the range [4, 128]. Note that we only perform hyper-parameter tuning on 100% (or fully observed) data, and use the same hyper-parameters for all experiments involving the dataset, such as different missing value probabilities. Table 7 captures the MissTSM parameters across datasets. Note that, for classification datasets (EMG, Epilepsy, Gesture), we use the same set of parameters with $q_{\text{dim}}=v_{\text{dim}}=k_{\text{dim}}=32$ and $n_{\text{head}}=16$. Table 8 presents the sensitivity analysis of the MissTSM-MAE model with respect to key hyperparameters (measured using MSE).
@@ C.2 Forecasting experiments
The models were trained with the MSE loss, using the Adam (Kingma, 2014) optimizer with a learning rate of 1e-3 during pre-training for 50 epochs and a learning r ate of 1e-4 during fine-tuning with an early stopping counter of 3 epochs. Batch size was set to 16. All the reported missing data experiment results are obtained over 5 trials (5 different masked versions). During fine-tuning for different Prediction lengths (96, 192, 336, 720), we used the same pre-trained encoder and added a linear layer at the top of the encoder.
@@ C.3 Classification experiments
The models were trained using the Adam (Kingma, 2014) optimizer, with MSE as the loss function during pre-training and Cross-Entropy loss during fine-tuning. During fine-tuning, we plugged a 64-D linear layer at the top of the pre-trained encoder. We pre-trained and fine-tuned for 100 epochs. For the GRU-D experiments, we use the available re-implementation of the code, with hidden size as 16. For the PhysioNet experiments, we follow the same evaluation setup as proposed in Du et al. (2023). The baselines imputation models use a simple RNN model as a classification model on top of the imputed data.
C.4 Discussion on Design Choices
We discuss some of the design choices behind the MissTSM layer.
Why cross-attention instead of a self-attention at each timestep? In our current formulation, we perform a masked cross-attention at every timestep using a single learnable query. While we agree that the conventional approach of applying self-attention can capture better cross-variate interactions, we specifically decided to learn a single query vector to reduce the computational cost of learning attention weights (we only need $N\times 1$ attention weights, instead of a full $N\times N$ matrix). While we empirically show that our simple attention architecture still captures valuable information across variates, our work can be easily extended in future works to model all variate-pairs interactions, e.g., by introducing a hierarchical or grouped aggregation mechanism that first models variate dependencies for a single time-step and then performs a global pooling across time-steps.
Why a single time-independent query? The goal of the MFAA layer is to learn a feature representation at each time-step using the observed variates while ignoring the missing variates. It does so by using a learnable query. The goal of learning this query vector can be equated to the task of extracting the best representation from the observed features at every timestep. Since this task can be assumed to be invariant of time (i.e., the same feature extractor can be used to learn representations of observed variates at any time-step), we consider learning a single time-independent query. Alternatively, if we used time-dependent queries, it would require learning $T$ separate query vectors, which would increase the number of learnable parameters. Moreover, by learning a different query for every time-step, the model would be forced to learn $T$ different rules, potentially making it susceptible to overfitting to artifacts at specific time indices.
Appendix D Complexity and Scalability Analysis
D.1 Theoretical Analysis of Computational Costs
We analyze the computational and memory complexity of the proposed MissTSM module. Let $L$ represent the context length, $N$ the number of variates (features), $d$, the embedding dimension, $h$, the number of heads, and $Q$, $K$, $V$ are the Query, Key and Value matrices
Table 9: Time, memory, and parameter complexity of MissTSM. Here, $L$ is the context length, $N$ the number of variates, $h$, number of heads and $d$ the embedding dimension. $d_{n}$ and $d_{l}$ are hidden and latent dimensions of mTAND, respectively. $G$ and $l_{inner}$ corresponds to the number of groups within the DMSA blocks and number of layers within each block in SAITS
Model Time Memory Params
MissTSM $\mathcal{O}(LNd^{2})$ $\mathcal{O}(LNd)$ $\mathcal{O}(d^{2}+Nd)$
mTAND $\mathcal{O}(L^{2}(d^{2}+h(N+d_{n}))$ $Ld_{l}d_{n})$ $\mathcal{O}(L(N+d_{n}+d_{l}+d)$ $hL^{2})$ $\mathcal{O}(d^{2}+hNd_{n}$ $hd_{n}^{2}+d_{n}d_{l})$
SimMTM $\mathcal{O}(NLdE(L+d))$ $\mathcal{O}(Ld+hL^{2})$ $\mathcal{O}(Ed^{2}+L^{2}d^{2})$
SAITS $\mathcal{O}(L(Nd+N^{2}+L)$ $Gl_{inner}L^{2}d)$ $\mathcal{O}(L(N+d)$ $Gl_{inner}(Ld+hL^{2}))$ $\mathcal{O}(l_{inner}d^{2}$ +N(d+N+L))+N(d+N+L))
Time Complexity The initial feature embedding (i.e. the Time-feature independent embedding TFI) processes each of the $N$ variates at each of the $L$ time steps, mapping them to a $d$-dimensional space. This operation has a time complexity of $O(LNd)$. The MFAA layer performs cross-attention across variates at each time-step, where a single query attends to $N$ embedded variates. The complexity is dominated by the linear projections to $Q$, and $K$ or $V$, $O(d^{2})$ and $O(Nd^{2})$ respectively. It results in a complexity of $O(LNd^{2})$. The final projection layer maps the $d$-dimensional output to the variate embedding space, adding $O(LdN)$ to the time complexity, overall resulting in a total time complexity of MissTSM as $O(L(Nd^{2}+dN))$, $O(LNd^{2})$
Parameter and Memory Complexity The main learnable parameters lie in the cross-attention layer (approximately $4d^{2}$ parameters from the $Q$, $K$, $V$ and the output projection layers), the $TFI$ embedding ($2Nd$ parameters from the weight and bias matrices), and the final linear layer ($dN$ parameters). Thus, the total parameter count is $d+4d^{2}+2Nd+2d$, yielding an overall parameter complexity of $O(d^{2}+Nd)$ - quadratic in $d$ and linear in NN
The largest intermediate tensors during the forward pass are the embedded features ($O(LNd)$) and the $Q$,$K$,$V$ tensors ($O(LNd)$), and the attention scores ($O(LN)$). Hence, the peak activation memory scales as $O(LNd)$, which is linear in $N$.
Table 9 summarizes the time, memory and parameter complexity of MissTSM, compared against three baselines - mTAND (IMTS method), SimMTM (MTS method) and SAITS (imputation method). The scalability of MissTSM is primarily governed by three factors - embedding dimension, number of variates and sequence length. In contrast, the baselines exhibit complexity that depends on multiple interacting components within each dimension (time, space, parameters), making their scalability less straightforward.
D.2 Empirical Analysis of Computational Costs
To evaluate the efficiency of different IMTS pipelines, we benchmark their training runtime, inference throughput, and peak GPU memory usage. Table 10 presents these metrics across models.
Table 10: Computational cost comparison between different pipelines and ours for IMTS modeling. We consider a classification task on the PhysioNet P19 dataset
Pipeline Training Time (avg. epoch, sec) $\downarrow$ Inference Throughput (samples/sec) $\uparrow$ Peak Active Tensor Memory (train, GB) $\downarrow$ # Learnable Params AUROC
$\mathbf{SAITS_{+MAE}}$ $13.277_{0.578}$ + $0.279_{0.163}$ = $13.556_{0.601}$ $438.135_{47.731}$ 0.74 $1,373,576+363,236$ = 1,736,812 77.40
mTAND $5.95_{0.117}$ $9593.36_{121.44}$ 1.6 142,370 82.9
Raindrop $17.651_{0.554}$ $964.75_{88.91}$ 1.4 1,947,804 87.6
$\mathbf{MissTSM_{+MAE}}$ $\textbf{0.322}_{0.077}$ $\textbf{19371.58}_{1315.92}$ 1.866 365,506 88.8
The training time is reported as the average per epoch over five trials. Inference throughput is computed as the total number of test samples divided by the total inference time. For two-stage IMTS pipeline, the total inference time is the sum across both the stages. We report the peak GPU memory allocated using PyTorch. For two-stage pipeline, we report the maximum GPU memory usage across both stages. All the experiments were conducted on Ubuntu 18.04, Python 3.9, PyTorch 1.12 (CUDA 12.8) and a single NVIDIA TITAN 24GB GPU, using a constant batch size of 64 for all models and with mixed precision disabled
[…]
D.3 MissTSM Scalability
We conducted an empirical study of runtime and memory usage as $N$ (number of variates) increases, to analyze the scalability of our $T\times N$ tokenization in the $TFI$ layer. We generated multiple synthetic datasets of approximately 56k observations with a 70% MCAR missing fraction, varying $N$ in 10,20,50,100,200,500,800,1000{10,20,50,100,200,500,800,1000}. For this analysis, we considered MissTSM integrated with an iTransformer backbone and compare its performance with a baseline approach of using SAITS imputation with iTransformer. We report the peak GPU memory and average time per epoch (Figure 12) as well as per forward pass (Figure 11). All the results are summarized in Table 11. In Figure 12, we compare the per epoch metrics between MissTSM+iTransformer and iTransformer model, and in Figure 11, we compare the metrics per forward pass between iTransformer and the MissTSM layer. The sequence length $T$ and the embedding dimension $d$ were held constant to 336 and 16 respectively, to isolate the effect of $N$. Note that the reported baseline results reflect only the runtime and memory requirements of the iTransformer model itself, excluding the additional computational cost associated with the SAITS imputation process, for ease of analysis.
Table 11: Scalability analysis with respect to NN.
N Forward Pass (per epoch) Forward + Backward Pass (per epoch)
Time (s) Peak GPU (GB) Time (s) Peak GPU (GB)
MissTSM iTrans MissTSM iTrans MissTSM+iTrans iTrans MissTSM+iTrans iTrans
10 2.095 1.784 0.080 0.064 22.416 13.495 0.271 0.177
20 2.954 1.911 0.149 0.114 27.312 16.275 0.410 0.236
50 6.004 3.163 0.359 0.275 49.903 18.514 0.847 0.427
100 10.227 12.487 0.711 0.576 96.618 48.387 2.764 1.925
200 20.140 24.188 1.410 1.309 174.575 69.040 3.313 1.636
500 52.917 62.014 3.515 4.536 456.296 185.875 9.182 5.034
800 83.468 113.758 5.613 9.309 1204.278 335.082 17.028 10.403
Figure 11: Time and peak GPU memory scaling w.r.t. NN, observed per forward pass
Figure 12: Time and peak GPU memory scaling w.r.t. NN, observed per epoch
From Figure 11, we observe that during the forward pass, the MissTSM layer scales almost linearly with $N$, consistent with the theoretical complexity in Table 9. However, Figure 12 shows that the runtime and memory for the combined forward and backward pass scale linearly only up to a certain point. Till $N=100$, iTransformer and MissTSM+iTransformer models scale almost in parallel, but diverges after that. Comparing the scaling curves in Figures 11 and 12, and the forward versus forward+backward metrics for iTransformer, we find that the backward pass dominates both runtime and peak GPU memory, leading to a sharp increase in runtime after $N>100$. On a single 24 GB NVIDIA RTX, training MissTSM becomes limiting at $N\approx 1000$. However, note that $N>100$ represents an extreme setup involving very large number of variates, while for common practical ranges of $N$ (e.g., $N<50$), the computational overhead of backpropagation with MissTSM remains minimal.
D.4 Imputation Time and Error Propagation - Case Study
We conduct a case study to examine the trade-offs between models for IMTS modeling in terms of computational complexity and efficiency, and to investigate potential correlations between imputation errors and downstream performance. We consider a classification task on the Epilepsy dataset that is 80% masked under MCAR. Spline and SAITS are the imputation techniques and SimMTM is the MTS model used. We report the total modeling time as the sum of imputation time and the time-series model training time. The experiments are conducted on NVIDIA TITAN 24 GB GPUs.
@@ E.3 Additional Forecasting Dataset
To further strengthen our understanding of the behaviors of regular IMTS models under the two-step pipeline, and to explore the potential of model-agnostic imputation-free approaches, we conduct forecasting experiments on another dataset, Solar-Energy (Lai et al., 2018), using the MCAR masking scheme over two missing data fractions - 60 and 70%. Figure 14 summarizes the results. As expected, the performance of traditional MTS methods degrades with increasing missingness, reflecting their reliance on separate imputation strategies. While IMTS methods provide a practical solution for partially observed time series, they tend to underperform compared to regular MTS models in long-range forecasting scenarios. This is where we believe the gap could potentially be addressed by imputation-free, model-agnostic approaches such as MissTSM.
Figure 14: Results on the Solar-Energy dataset with MCAR masking across four horizons at 60% and 70% masking fractions. Lower MSE is better
E.4 Analyzing the Impact of Forecast Horizon